# Revisiting Experience Replayable Conditions

### v1

@@ 4.1 Experience discriminator
[…]
Figure 3: Two stabilization tricks
The instability factors induced by the above triplet loss have to be suppressed to satisfy ERC. To this end, two stabilization tricks, counteraction and mining, are heuristically designed (see Fig. 3). As a common module for them, experience discriminator, is first introduced.
@@
where $\sigma(\cdot)\in[0,1]$ is the sigmoid function. Note that $d$ should be equal to or less than $0.5$ since $(s,a)$ is actually generated from the behavior policy with likelihood $b$.
In addition, robust judgements should be considered w.r.t. the stochasticity of actions and the non-stationarity of policies. For this purpose, $D:\mathcal{S}\mapsto[0,1]$, which marginalizes $d$ by $a$ and has only $s$ as input, is defined as a learnable model. As $D$ corresponds to the probability parameter of Bernoulli distribution, it can be optimized through minimization of its negative log-likelihood.
@@ 4.2 Counteraction of deviations from non-optimal policies
where, $\lambda\geq 0$ denotes the gain for imbalancing the positive and negative terms  [^5] The same is true if the second term is multiplied by the gain $\lambda\in(0,1)$..
Now, how this extension works in practice is shown. As in the original minimization problem, the gradient for the added regularization is derived as below.
@@
That is, it takes advantage of the fact that the higher the misidentification rate of $D$ means $\pi\sim b$, reducing the risk of hard negative (see Fig. 3). Since $d$ has the computational graph w.r.t. $\pi$, the following loss function can be given as the counteraction trick for regularizing $\pi$ to $b$.
@@
where, $\omega\geq 0$ denotes the gain, which is designed below. One can find that the term related to $D$ also behaves as a gain and is larger when $D\ll 1$ (i.e. no misidentification). In addition, if $d$ is clipped to $0.5$ (i.e. $\pi\simeq b$ holds), the gradient w.r.t. $\pi$ is zero. Note that, in the actual implementation, the gradient reversal layer [12] is useful to lump eqs. (16) and (20) together.
As for $\omega$, $d\ll 1$ has the small gradient due to the sigmoid function, not like $D\ll 1$. To compensate it and improve convergence performance to $\pi\to b$ even in such a case, $\omega$ is designed in the manner of PI control, referring to the literature [45].
@@ 4.3 Mining of indistinguishable experiences
Second trick, so-called the mining, is proposed to mitigate the effect of distribution shift mainly. In previous studies, semi-hard triplets, which satisfy $d(x,x^{+})+m<d(x,x^{-})$, are considered useful [41]. Since the optimization problem in this study sets $m=0$, $\mathrm{KL}(\pi\mid\pi^{+})<\mathrm{KL}(\pi\mid\pi^{-})$ seems better. However, this is also regarded to be hard triplets, which would induce selection bias. Of course, $\mathrm{KL}(\pi\mid\pi^{+})>\mathrm{KL}(\pi\mid\pi^{-})$ is still not suitable for learning because of the hard negative relationship described above. Hence, the triplets with $\mathrm{KL}(\pi\mid\pi^{+})\simeq\mathrm{KL}(\pi\mid\pi^{-})$ are desired to be mined.
Specifically, it takes advantage of the fact that, in this study, the anchor data is determined by the current policy $\pi$, and the positive and negative data, $\pi^{+}$ and $\pi^{-}$, are located around the behavior policy $b$. In other words, the indistinguishable empirical data with $\pi\simeq b$ is likely to achieve the desired relationship (see Fig. 3). The mining trick is therefore given as the following stochastic dropout [44].
@@ 5.1 Overview
Here, it is verified that ERC is satisfied by the two stabilization tricks. To do so, A2C [33], which is the on-policy algorithm and generally considered inapplicable for ER, is employed as a baseline (see Appendix A for detailed settings). Learning is conducted only with the empirical data replayed from ER at the end of each episode (without online learning).
The following three steps are performed for the step-by-step verification. With them, it is shown that alleviating the instability factors hidden in triplet loss is effective not only to satisfy ERC, but also to learn with comparable sample efficiency to SAC [14].
@@
    Finally, A2C with the proposed stabilization tricks is evaluated on two tasks of relatively large problem scale in dm_control [48] in comparison to SAC, which is the latest off-policy algorithm.
Note that training for each task and condition is conducted 12 times with different random seeds in order to evaluate the performance with their statistics.
@@ 5.2 Effective range of hyperparameters
Figure 4: Grid search of $(\eta_{C},\eta_{M})$ on DoublePendulum
@@
Based on the two trends, and as a natural setting, $(\eta_{C},\eta_{M})$ are decided to be $(0.5,2.0)$. The learning performance at this time is also shown in Fig. 4 (the black bar), and is higher than the rough grid search results (i.e. $\sim 416$ while the others did not exceed $400$).
@@ 5.3 Ablation tests
[…]
Figure 5: Returns of ablation tests
[…]
Figure 6: $\omega(s,a,b)$ in the counteraction trick
[…]
Figure 7: $p(M\mid s,a,b)$ in the mining trick
@@
Now let’s see why both stabilization tricks are required. To this end, the internal parameters for the respective stabilization tricks, $\omega(s,a,b)$ and $p(M\mid s,a,b)$, are depicted in Figs. 6 and 7. One can find that the cases only with one of the stabilization tricks immediately saturated the respective internal parameters. That is, if using the counteraction trick only, its regularization to $\pi\to b$ is not enough for satisfying ERC; and if using the mining trick only, its selection is too strict for learning the optimal policy. On the other hand, by using both, the internal parameters converged to task-specific values without saturation. That is why the two stabilization tricks need to be used together to increase the amount of replayable empirical data to reach a level where the optimal policy can be learned while still satisfying ERC.
@@ 5.4 Performance comparison
Figure 8: Test results on practical tasks
@@
The test results after learning are depicted in Fig. 8. As can be seen, while SAC solved Quadruped at a high level, it did not solve Swimmer consistently. On the other hand, ERC performed unstably on Quadruped, but showed high performance on Swimmer. Thus, it can be said that ERC has a learning performance comparable to that of SAC, although they have different strengths in different tasks.
As a remark, the difference in performance between ERC and SAC might be attributed to the difference in the policy gradients used. Actually, ERC learns the policy with likelihood-ratio gradients, while SAC does so with reparameterization gradients [37]. The former obtains the gradients over the entire (multi-dimensional) policy, which cannot optimize each component of actions independently. As a result, this is useful to learn tasks that require synchonization of actions. The latter, on the other hand, obtains the gradients for each component of actions, and thus can quickly learn tasks that do not require precise synchronization. These learning characteristics correspond to the two types of tasks solved in this study. Hence, it is suggested that SAC performed well for Quadruped, where synchronization at each leg is sufficient, and ERC performed well for Swimmer, where synchronization at all joints is needed. Note that policy improvements by mixing them have been proposed, which would be expected to be useful [13].
@@ 6.1 Relations to conventional algorithms
As demonstrated in the above section, ERC hypothesized in this study can be achieved by limiting the replayed empirical data to ones acceptable by the applied RL algorithm. At the same time, the improvement in learning efficiency due to ER can be increased by expanding the acceptable set of empirical data. These were achieved by the two stabilization tricks proposed, with $\pi\simeq b$ being the key to both. This corresponds to the concept of on-policyness, where the empirical data contained in the replay buffer can be regarded as generated from the current policy $\pi$ [10]. In other words, the proposed stabilization tricks aimed to stabilize the triplet loss and consequently increased the on-policyness. Conversely, the usefulness of increasing the on-policyness in previous studies [34, 43] can be explained in terms of mitigating the instabilization factors hidden in the triplet loss.
@@
Actually, ER can be applied to make A2C off-policy [8]. To this end, the most important trick is to change the sampler by importance sampling, as described above. This trick involves weighting the original loss function by $\pi/b$, which is unstable when $b\simeq 0$ (i.e. rare actions are sampled). Several heuristics have been proposed to resolve this instability [52, 57, 9]. On the other hand, the proposed stabilization tricks selectively replay the empirical data with $\pi\simeq b$ while increasing the number of acceptable ones. As a result, the weight in this important sampling can be ignored as $\pi/b\simeq 1$. That is why A2C with the proposed stabilization tricks, although not explicitly off-policy, made ER available as if they were off-policy.
@@
As a remark, the report that PPO (and its variants), which is on-policy [42], could utilize ER empirically [26] is related to the above considerations and the proposed stabilization tricks. That is, PPO applies regularization aiming at $\pi\simeq b$, and if $\pi$ deviates from $b$, it clips the weight of the importance sampling, resulting in excluding that data. Thus, PPO can increase the on-policyness of empirical data and exclude from replays that are unacceptable, so ER is actually applicable.
@@
Finally, another interpretation of why DDPG [30], SAC [14], and TD3 [11], which are considered off-policy, can stably learn the optimal policy using ER. These algorithms optimize the current policy $\pi$ by using reparameterization gradients through the actions generated from $\pi$ fed into the action value function $Q$. As mentioned in the SAC paper, this can be derived from the minimization of KL divergence between $\pi$ and the optimal policy based on $Q$ only. In other words, unlike algorithms that use likelihood-ratio gradients such as A2C, they do not deal with the triplet loss. Therefore, it can be considered that ER is applicable since there is no instabilization factors induced by the triplet loss in the first place.
@@ 6.2 Limitations of stabiliation tricks
For ERC holding, this paper proposed the two stabilization tricks, the contributions of which were evaluated in the above section. However, these naturally leave room for improvements. First, using them requires a discriminator $D$ learned through eq. (16), which needs an extra computational cost. While it is possible to implement the two tricks using only $d$ in eq. (15) without $D$, $d$ is an unstable variable, so a lightweight stabilizer will be needed to replace $D$.
It is also important to note that the decision on $D$ and $d$ is based on the likelihood $\pi$ and $b$. That is, when the policy deals with a high-dimensional action space, even a small deviation in action may be judged as $\pi\neq b$ severely. In fact, when the RL benchmark with musculoskeletal model, i.e. myosuite [4], was tested with $|\mathcal{A}|=39$, $p(M\mid s,a,b)$ often converged to one depending on the conditions. This means that most of the empirical data could not be replayed. To avoid this excessive exclusion of empirical data, it would be useful to either determine the final exclusion of empirical data by summarizing the respective judgements on individual action dimension separately; or to optimize the policy by masking the inappropriate action dimension only.
In addition, the counteraction alone did not satisfy ERC, as indicated in the ablation tests. This is probably due to saturation of the regularization gain $\omega$ (see Fig. 6). Although a non-saturated gain made learning unstable empirically, it would be better to relax the saturation to some extent. Alternatively, an auto-tuning trick, which is possible by once interpreting the regularization as the constraint [15, 27], is considered to be useful.
@@ 7 Conclusion
This study reconsidered the factors that cause whether ER is applicable to RL algorithms. To this end, the instability factors that ER might induce especifally in on-policy algorithms were first revealed. In other words, it was found that the policy gradient algorithms can be regarded as the minimization of triplet loss in metric learning, inheriting its instability factors. To alleviate them, the two stabilization tricks, the counteraction and mining, were proposed as countermeasures attached to arbitrary RL algorithms. The counteraction and mining are responsible for i) expanding the set of acceptable empirical data for each RL algorithm and ii) excluding empirical data outside the set, respectively. Through multiple simulations, ERC indeed holded by using these two stabilization tricks. Furthermore, the standard on-policy algorithm with them achieved the learning performance comparable to the state-of-the-art off-policy algorithm.
As described in the discussion, the two stabilization tricks proposed to satisfy ERC leave some room for improvements. In particular, since the hypothesis formulated in this study is now deeply related to the on-policyness of the empirical data in the replay buffer, they would be improved based on this perspective. Afterwards, other RL algorithms will be integrated with the stabilization tricks for further investigations of ERC.

### v2

@@ 4.1 Experience discriminator
(a) Counteraction
(b) Mining
Figure 3: Two stabilization tricks: According to the judgements of experience discriminator, (a) counteraction applies a regularization to $\pi$ to increase the misidentification rate to $\pi\simeq b$; and (b) mining masks empirical data judged to be $\pi\neq b$ and excludes them from replay/training.
The instability factors induced by the above triplet loss have to be suppressed to satisfy ERC. To this end, two stabilization tricks, counteraction and mining, are heuristically designed (see Fig. 3). As a common module for them, experience discriminator, is first introduced. In other words, it estimates whether the empirical data of ER is suitable for triplet construction.
@@
where $\sigma(\cdot)\in[0,1]$ is the sigmoid function. Note that $d$ should be equal to or less than $0.5$ since $(s,a)$ is actually generated from the behavior policy with its likelihood $b$.
In addition, robust judgements should be considered w.r.t. the stochasticity of actions and the non-stationarity of policies. For this purpose, $D:\mathcal{S}\mapsto[0,1]$, which marginalizes $d$ by $a$ and has only $s$ as input, is defined as a learnable model. As $D$ corresponds to the probability parameter of Bernoulli distribution, it can be optimized by minimizing its negative log-likelihood.
@@ 4.2 Counteraction of deviations from non-optimal policies
where $\lambda\geq 0$ denotes the gain for imbalancing the positive and negative terms  [^6] The same is true if the second term is multiplied by the gain $\lambda\in(0,1)$..
Now, how this extension works in theory is shown. As in the original minimization problem, the gradient for the added regularization is derived as below.
@@
That is, it takes advantage of the fact that the higher the misidentification rate of $D$ means $\pi\sim b$, reducing the risk of hard negative (see Fig. 3(a)). Since $d$ has the computational graph w.r.t. $\pi$, the following loss function can be given as the counteraction trick for regularizing $\pi$ to $b$.
@@
where $\omega\geq 0$ denotes the gain, which is designed below. One can find that the term related to $D$ also behaves as a gain and is larger when $D\ll 1$ (i.e. no misidentification). In addition, if $d$ is clipped to $0.5$ (i.e. $\pi\simeq b$ holds), the gradient w.r.t. $\pi$ is zero. Note that, in the actual implementation, the gradient reversal layer [15] is useful to lump eqs. (16) and (20) together.
As for $\omega$, $d\ll 1$ has the small gradient due to the sigmoid function, not like $D\ll 1$. To compensate it and converge to $\pi\to b$ faster even in such a case, $\omega$ is designed in the manner of PI control, referring to the literature [50].
@@ 4.3 Mining of indistinguishable experiences
Second trick, so-called the mining, is proposed to mitigate the effect of distribution shift mainly. In the previous studies, semi-hard triplets, in which $d(x,x^{+})+m<d(x,x^{-})$ holds, are considered useful [45]. Since the optimization problem in this study sets $m=0$, $\mathrm{KL}(\pi\mid\pi^{+})<\mathrm{KL}(\pi\mid\pi^{-})$ seems better. However, this is also regarded to be hard triplets, which would induce the selection bias. Of course, $\mathrm{KL}(\pi\mid\pi^{+})>\mathrm{KL}(\pi\mid\pi^{-})$ is still not suitable for learning because of the hard negative relationship described above. Hence, the triplets with $\mathrm{KL}(\pi\mid\pi^{+})\simeq\mathrm{KL}(\pi\mid\pi^{-})$ are desired to be mined.
Specifically, it takes advantage of the fact that, in this study, the anchor data is determined by the current policy $\pi$, and the positive and negative data, $\pi^{+}$ and $\pi^{-}$, are located around the behavior policy $b$. In other words, the indistinguishable empirical data with $\pi\simeq b$ is likely to achieve the desired relationship (see Fig. 3(b)). The mining trick is therefore given as the following stochastic dropout [49].
@@ 5.1 Overview
Here, it is verified that ERC holds by the two stabilization tricks. To do so, A2C [37], which is the on-policy algorithm and generally considered inapplicable for ER, is employed as a baseline (see Appendix A for detailed settings). Learning is conducted only with the empirical data replayed from ER at the end of each episode (without online learning).
The following three steps are performed for the step-by-step verification. With them, it is shown that alleviating the instability factors hidden in triplet loss is effective not only to satisfy ERC, but also to learn the optimal policy with comparable sample efficiency to SAC [17].
@@
    Finally, the A2C with the proposed stabilization tricks is evaluated on two tasks of relatively large problem scale in dm_control [54] in comparison to SAC, which is the latest off-policy algorithm.
Note that training for each task and condition is conducted 12 times with different random seeds in order to evaluate the performance with their statistics. In the above, the stabilization tricks demonstrate that ER can be applied to A2C with the excellent learning performance. In addition, the application of the two stabilization tricks to the other on-policy algorithm, i.e. PPO [47], and their quantitative contribution to the learning performance are investigated in Appendix B.
@@ 5.2 Effective range of hyperparameters
Figure 4: Grid search of $(\eta_{C},\eta_{M})$ on DoublePendulum: $\eta_{C,M}=\{0.1,0.5,1.0,5.0,10.0\}$ are searched roughly in terms of return from the environment, and then $(\eta_{C},\eta_{M})$ used later are decided to be $(0.5,2.0)$ referring to the grid search results.
@@
Based on the above two trends, and as a natural setting, $(\eta_{C},\eta_{M})$ are decided to be $(0.5,2.0)$ for the remaining experiments. The learning performance at this time is also shown in Fig. 4 (the black bar), and is higher than the rough grid search results (i.e. $\sim 416$ while the others did not exceed $400$).
@@ 5.3 Ablation tests
(a) Reacher
(b) Hopper
(c) HalfCheetah
Figure 5: Returns of ablation tests: The addition of the two stabilization tricks enabled the learning with an on-policy algorithm, A2C, combined with ER.
(a) Reacher
(b) Hopper
(c) HalfCheetah
Figure 6: $\omega(s,a,b)$ in the counteraction trick: While the counteraction trick alone saturated the regularization gain and made it difficult to perform $\pi\simeq b$, the two tricks made the regularization work properly without saturation.
(a) Reacher
(b) Hopper
(c) HalfCheetah
Figure 7: $p(M\mid s,a,b)$ in the mining trick: While the mining trick alone saturated the dropout rate and made it unavailable for replaying enough empirical data, the two tricks made the selection work properly without saturation.
@@
Now let’s see why both of the stabilization tricks are required. To this end, the internal parameters for the respective stabilization tricks, $\omega(s,a,b)$ and $p(M\mid s,a,b)$, are depicted in Figs. 6 and 7. One can find that the cases only with one of the stabilization tricks saturated the respective internal parameters immediately. That is, if using the counteraction trick only, its regularization to $\pi\to b$ is not enough for satisfying ERC; and if using the mining trick only, its selection is too strict for learning the optimal policy. In contrast, by using both, the internal parameters converged to task-specific values without saturation. That is why the two stabilization tricks need to be employed together to increase the amount of replayable empirical data to reach a level where the optimal policy can be learned while still satisfying ERC.
@@ 5.4 Performance comparison
(a) Quadruped
(b) Swimmer
(c) Returns after learning
Figure 8: Returns on practical tasks: From (a) and (b), it can be found that SAC has faster learning speed, while ERC has more stable learning ability, resulting in (c) learning performance of ERC comparable to SAC.
@@
The learning curves and test results after learning are depicted in Fig. 8. As can be seen, while SAC solved Quadruped at a high level, it did not solve Swimmer consistently. On the other hand, ERC sometimes delayed convergence on Quadruped, but showed high performance on Swimmer. In addition, the trends of the learning curves suggested that SAC learns faster, while ERC learns more stably. Thus, it can be said that ERC has a learning performance comparable to that of SAC, although they have different strengths in different tasks.
As a remark, the difference in performance between ERC and SAC might be attributed to the difference in the policy gradients used. Actually, ERC learns the policy with likelihood-ratio gradients, while SAC does so with reparameterization gradients [41]. The former obtains the gradients over the entire (multi-dimensional) policy, which cannot optimize each component of actions independently. As a result, this is useful to learn tasks that require synchonization of actions. The latter, on the other hand, obtains the gradients for each component of actions, and thus can quickly learn tasks that do not require precise synchronization. These learning characteristics correspond to the two types of tasks solved in this study. Hence, it is suggested that SAC performed well for Quadruped, where synchronization at each leg is sufficient, and ERC performed well for Swimmer, where synchronization at all joints is needed. Note that policy improvements by mixing them have been proposed, which would be expected to be useful [16]. The proposed stabilization tricks allow for the policy improvements combined with ER regardless of which gradient is used. That would facilitate the development of such hybrid algorithms and improve learning performance in a complementary manner.
@@ 6.1 Relations to conventional algorithms
As demonstrated in the above section, ERC hypothesized in this study can be achieved by limiting the replayed empirical data to ones acceptable by the applied RL algorithm. At the same time, the improvement in learning efficiency due to ER can be increased by expanding the acceptable set of empirical data. These were achieved by the two stabilization tricks proposed, with $\pi\simeq b$ being the key to both. This key corresponds to the concept of on-policyness, where the empirical data contained in the replay buffer can be regarded as generated from the current policy $\pi$ [13]. In other words, the stabilization tricks aimed to stabilize the triplet loss and consequently increased the on-policyness. Conversely, the usefulness of increasing the on-policyness in the previous studies [38, 48] can be explained in terms of mitigating the instability factors hidden in the triplet loss.
@@
Actually, ER can be applied to make A2C off-policy [11]. To this end, the most important trick is to change the sampler by importance sampling, as described above. This trick involves weighting the original loss function by $\pi/b$, which is unstable when $b\simeq 0$ (i.e. rare actions are sampled). Several heuristics have been proposed to resolve this instability [58, 64, 12]. Specifically, the most famous off-policy actor-critic algorithm with ER, ACER [58], modifies the update rules of not only the policy function but also the value function, while the proposed tricks indicated no need to modify the value function. Although Geoff-PAC [64] and P3O [12] only modify the policy improvements, the former is tedious and lacks the intuitive explanatory capability of ERC, while the latter is inappropriate for confirming ERC because it uses the freshly collected empirical data for learning separately from those sampled from ER.
Thus, previous studies have only focused on off-policyness and improved learning performance of the algorithms, without any discussion why ER is applicable. On the other hand, this study explored ERC and implicitly obtained utility similar to making the target algorithm off-policy. Indeed, the proposed stabilization tricks selectively replay the empirical data with $\pi\simeq b$ while increasing the number of acceptable ones. As a result, the weight in this important sampling can be ignored as $\pi/b\simeq 1$. That is why the A2C with the proposed stabilization tricks, although not explicitly off-policy, made ER available as if they were off-policy.
@@
As a remark, the report that PPO (and its variants), which is regarded to be on-policy [47], could utilize ER empirically [28] is related to the above considerations and the proposed stabilization tricks. That is, PPO applies regularization aiming at $\pi\simeq b$, and if $\pi$ deviates from $b$, it clips the weight of the importance sampling, resulting in excluding that data. Thus, PPO can increase the on-policyness of empirical data and exclude from replays that are unacceptable, so ER is actually applicable (see the empirical results in Appendix B).
@@
Finally, another interpretation of why DDPG [33], SAC [17], and TD3 [14], which are considered off-policy, can stably learn the optimal policy using ER. These algorithms optimize the current policy $\pi$ by using reparameterization gradients through the actions generated from $\pi$ fed into the action value function $Q$. As mentioned in the SAC paper, this can be derived from the minimization of KL divergence between $\pi$ and the optimal policy based on $Q$ only. In other words, unlike algorithms that use likelihood-ratio gradients such as A2C, they do not deal with the triplet loss. Therefore, it can be considered that ER is applicable since there is no instability factors induced by the triplet loss in the first place.
@@ 6.2 Limitations of stabiliation tricks
For ERC holding, this paper proposed the two stabilization tricks, the contributions of which were evaluated in the above section. However, these naturally leave room for improvements probably because they have been designed heuristically. First, using them requires a discriminator $D$ learned through eq. (16), which needs an extra computational cost. While it is possible to implement the two stabilization tricks using only $d$ in eq. (15) without $D$, $d$ is an unstable variable, so a lightweight stabilizer will be needed to replace $D$.
It is also important to note that the decision on $D$ and $d$ is based on the likelihood $\pi$ and $b$. That is, when the policy deals with a high-dimensional action space, even a small deviation in action may be judged as $\pi\neq b$ severely. In fact, when the RL benchmark with musculoskeletal model, i.e. myosuite [6], was tested with $|\mathcal{A}|=39$, $p(M\mid s,a,b)$ often converged to one depending on the conditions. This means that the empirical data were rarely replayed. To avoid this excessive exclusion of empirical data, it would be useful to either determine the final exclusion of empirical data by summarizing the respective judgements on individual action dimension separately; or to optimize the policy by masking the inappropriate action dimension only.
In addition, the counteraction alone did not satisfy ERC, as indicated in the ablation tests. This is probably due to saturation of the regularization gain $\omega$ (see Fig. 6). Although a non-saturated gain made learning unstable empirically, it would be better to relax the saturation to some extent. Alternatively, an auto-tuning trick, which is possible by once interpreting the regularization as the corresponding constraint [18, 29], is considered to be useful.
@@ 7 Conclusion
This study reconsidered the factors that determine whether ER is applicable to RL algorithms. To this end, the instability factors that ER might induce especially in on-policy algorithms were first revealed. In other words, it was found that the policy gradient algorithms can be regarded as the minimization of triplet loss in metric learning, inheriting its instability factors. To alleviate them, the two stabilization tricks, the counteraction and mining, were proposed as countermeasures attached to arbitrary RL algorithms. The counteraction and mining are responsible for i) expanding the set of acceptable empirical data for each RL algorithm and ii) excluding empirical data outside the set, respectively. Through multiple simulations, ERC indeed holded by using these two stabilization tricks. Furthermore, the standard on-policy algorithm with them achieved the learning performance comparable to the state-of-the-art off-policy algorithm.
As described in the discussion, the two stabilization tricks proposed to satisfy ERC leave some room for improvements. In particular, since the hypotheses formulated in this study is now deeply related to the on-policyness of the empirical data in the replay buffer, they would be improved based on this perspective. Afterwards, other RL algorithms will be integrated with the stabilization tricks for further investigations of ERC. On the other hand, although the simplest ER method was tested for simplicity since the proposed tricks allow for arbitrary ER methods, it would be interesting to consider more sophisticated methods [35, 59] and methods modified to suit the problem [9] or algorithm [1]. In particular, by following the latter direction and developing an ER method suitable for on-policy RL algorithms, a significant performance improvement might be expected.